# Fractional $p$-Laplace Systems with critical Hardy Nonlinearities: Existence and Multiplicity

### v1

@@ 2. Existence and Non-existence of ground state solutions
$0<\frac{\delta_{1}}{2}=r_{n}^{d-sp}\fint_{B(y_{n},r_{n})}(|{u}_{n}|^{p}+|{v}_{n}|^{p})\,{\rm d}x\leq Cr_{n}^{-sp}\int_{\Omega}(|{u}_{n}|^{p}+|{v}_{n}|^{p})\,{\rm d}x.$
If $r_{n}\to\infty$, then using the fact that $\{u_{n}\},\{v_{n}\}$ are bounded in $L^{p}(\Omega)$, we get a contradiction. Thus, $\{r_{n}\}$ is a bounded sequence. By construction, this implies that $\{y_{n}\}$ is bounded in ${\mathbb{R}}^{d}$. Suppose, $r_{n}\to r_{0}>0$ and $y_{n}\to y\in{\mathbb{R}}^{d}$. Notice that $\tilde{u}_{n}$ and $\tilde{v}_{n}$ are supported on $\frac{\Omega-y_{n}}{r_{n}}$ and as $n\to\infty$,
@@
Since $(\tilde{u}_{n},\tilde{v}_{n})\to(\tilde{u},\tilde{v})$ in ${\mathcal{W}}(\mathbb{R}^{d})$, $(\tilde{u},\tilde{v})$ satisfies,
@@
Thus, $(\tilde{u},\tilde{v})$ minimizes $S_{\alpha,\beta}$ on a bounded set, which is absurd. Hence $r_{n}\to 0$ and since $y_{n}\in\{x\in{\mathbb{R}}^{d}:d(x,\overline{\Omega})\leq r_{n}\}$, we must have $y\in\overline{\Omega}$. ∎
@@ 3. Concentration compactness principle associated with the homogeneous system
Let $0\leq m\leq sp<d$ and $\alpha,\beta>1$ be such that $\alpha+\beta=p^{*}_{s}(m)$. Let $\{(u_{n},v_{n})\}$ weakly converges to ${(u,v)}$ in $\mathcal{W}$. Then up to a subsequence, the following identities hold:
@@
Let $\Omega\subset{\mathbb{R}}^{d}$ be a bounded open set containing 00. Let $0\leq m\leq sp$ and $\alpha,\beta>1$ be such that $\alpha+\beta=p^{*}_{s}(m)$. Let $\{u_{n}\},\{v_{n}\}$ be sequences in ${W_{0}^{s,p}(\Omega)}$ such that $u_{n}\rightharpoonup u$ and $v_{n}\rightharpoonup v$ in ${W_{0}^{s,p}(\Omega)}$. Then there exist measures $\sigma_{1},\sigma_{2},\nu_{1},\nu_{2},\nu$ and at most countable sets $\{x_{j}\}_{j\in\mathcal{J}_{1}}$, $\{x_{j}\}_{j\in\mathcal{J}_{2}}$, $\{x_{j}\}_{j\in\mathcal{J}_{3}}$, such that up to subsequences,
@@
which implies that $\{x_{j}\}_{j\in\mathcal{J}_{3}}=\{0\}$. Suppose there exists some $x_{j}\in\mathcal{J}_{3}$ with $x_{j}\neq 0$ and $x_{j}\in\text{supp}(\nu)$, then choosing $\phi\in{{\mathcal{C}}_{c}^{\infty}}(\Omega)$ and $\phi\equiv 1$ in a neighbourhood of $x_{j}$ with support does not contain 00, we get a contradiction repeating our previous calculation. Next, for $u\neq 0$ and $v\neq 0$, taking $\tilde{u}_{n}=u_{n}-u$ and $\tilde{v}_{n}=v_{n}-v$, we immediately get
@@ 4. Existence of a nontrivial solution with least energy
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
By [27, Theorem 1.1], we choose a positive radially symmetric decreasing minimizer $U_{m}=U_{m}(r)$ for $S_{\alpha+\beta}$ which weakly solves
@@
For $\varepsilon>0$, the function $U_{m,\varepsilon}(x)=\varepsilon^{-\frac{N-ps}{p}}U_{m}\left(\frac{x}{\varepsilon}\right)$ is also a minimizer of $S_{\alpha+\beta}$ satisfying (4.11). Now, for $\delta>0$, we define the radially symmetric non-decreasing function
@@
$\begin{cases}r>p^{*}_{s}-p^{\prime},&\text{ if }d<p^{2}s;\\ r\geq p,&\text{ if }d\geq p^{2}s.\end{cases}$ (4.13)
@@
$I\left(tu_{0},tv_{0}\right):=g_{\varepsilon}(t)$ $=\frac{1}{p}\left(\left[tu_{0}\right]_{s,p}^{p}+\left[tv_{0}\right]_{s,p}^{p}\right)-\frac{\gamma}{p_{s}^{*}(m)}\int_{\Omega}\frac{\left(tu_{0}\right)^{\alpha}\left(tv_{0}\right)^{\beta}}{|x|^{m}}\,{\rm d}x$
@@
$\quad\quad-\frac{\eta t^{r}}{r}\left(\alpha^{\frac{r}{p}}+\beta^{\frac{r}{p}}\right)\int_{\Omega}u_{m,\alpha,1}^{r}\,{\rm d}x$
@@
$l_{S}:=\displaystyle\inf_{(u,v)\in S}I(u,v)=I(\tilde{u},\tilde{v}),$
@@
Then ($\mathcal{S}$) admits a nontrivial solution $(\tilde{u},\tilde{v})$ with the least energy provided
@@
By Lemmata 4.2, 4.4 and 4.5, and applying the mountain pass theorem [3, Theorem 2.1], there exists a critical point $(\tilde{u},\tilde{v})\in{\mathcal{W}}$ for the functional $I$ which implies $(\tilde{u},\tilde{v})$ is a solution for ($\mathcal{S}$). Further, $I(\tilde{u},\tilde{v})=c_{1}>0$ implies that $(\tilde{u},\tilde{v})$ is a nontrivial solution. Next, we prove that $(\tilde{u},\tilde{v})$ has the least energy, i.e., $I(\tilde{u},\tilde{v})=l_{S}$. By the definition of $l_{\mathcal{N}}$ and $l_{S}$, we have
$l_{\mathcal{N}}\leq l_{S}\leq c_{1}=I(\tilde{u},\tilde{v}).$ (4.18)
@@
$[u]_{s,p}^{p}+[v]_{s,p}^{p}=t^{p_{s}^{\ast}(m)-p}\gamma\int_{\Omega}\frac{\left|u\right|^{\alpha}\left|v\right|^{\beta}}{|x|^{m}}\,{\rm d}x+t^{r-p}\eta\int_{\Omega}\left(\left|u\right|^{r}+\left|v\right|^{r}\right)\,{\rm d}x.$
@@
$l_{\mathcal{N}}=l_{S}=c_{1}=I(\tilde{u},\tilde{v}),$
which proves that $(\tilde{u},\tilde{v})$ has the least energy. ∎
Next, we show that $(\tilde{u},\tilde{v})$ is non semi-trivial. Take $\tilde{\eta}\geq 0$. To show the dependency on $\tilde{\eta}$, we denote $I,{\mathcal{N}},\tilde{u},\tilde{v}$ by $I_{\tilde{\eta}},{\mathcal{N}}_{\tilde{\eta}},u_{\tilde{\eta}},v_{\tilde{\eta}}$ respectively. In view of this,
$\theta(\tilde{\eta}):=\inf_{{\mathcal{N}}_{\tilde{\eta}}}I_{\tilde{\eta}}=I_{\tilde{\eta}}(u_{\tilde{\eta}},v_{\tilde{\eta}})<\tilde{c}.$
@@
Observe that $I_{+}\in C^{1}({\mathcal{W}},{\mathbb{R}})$ and every critical point of $I_{+}$ corresponds to a non-negative weak solution of ($\mathcal{S}$). We remark that Lemmas 4.24.5 hold for $I_{+}$ as well. As a consequence, we get a non-negative, non semi-trivial solution $(\hat{u},\hat{v})$ of ($\mathcal{S}$), provided $\gamma,\eta$ and $r$ as in Theorem 1.4. Further, taking $(\hat{u},0)$ and $(0,\hat{v})$ as test functions in ${\mathcal{W}}$ and using the fact that $(\hat{u},\hat{v})$ weakly solves ($\mathcal{S}$), we see that the following hold weakly:
@@ 5. Existence of category many nontrivial solutions
$\eta<\min\left\{\frac{\lambda_{1}}{2},\frac{S_{\alpha+\beta}}{c_{0}}\right\}\text{ when }r=p,\text{ and }\begin{cases}r>p^{*}_{s}-p^{\prime},&\text{ if }d<p^{2}s;\\ r\geq p,&\text{ if }d\geq p^{2}s.\end{cases}$
We define $\psi_{\eta}:{\mathcal{W}}\to{\mathbb{R}}$ by
@@
[…]
@@
Now our aim is to show that for $\eta$ small enough, $\text{cat}_{I_{\eta}^{\theta(\eta)}}(I_{\eta}^{\theta(\eta)})$ is bigger than $\text{cat}_{\Omega}(\Omega).$ For $\tau>0$, we define
@@
There exists $\eta_{\ast\ast}>0$ such that for $\eta\in(0,\eta_{\ast\ast})$, $\Phi_{\eta}(u,v)\in\Omega_{\tau}^{+}$ for every $(u,v)\in I_{\eta}^{\theta(\eta)}$.
@@
Arguing by contradiction, suppose there exists $\eta_{n}>0$, $\eta_{n}\to 0$ and $(u_{n},v_{n})\in I_{\eta_{n}}^{\theta(\eta_{n})}$ such that $\Phi_{\eta_{n}}(u_{n},v_{n})\not\in\Omega_{\tau}^{+}$. Then
@@
$\frac{1}{p}\|(u_{n},v_{n})\|_{{\mathcal{W}}}^{p}-\frac{\gamma}{p_{s}^{*}}\int_{\Omega}{|u_{n}|^{\alpha}|v_{n}|^{\beta}}\,{\rm d}x-\frac{\eta_{n}}{r}\int_{\Omega}(|u_{n}|^{r}+|v_{n}|^{r})\,{\rm d}x\leq{\theta(\eta_{n})}<\tilde{c}.$ (5.4)
@@
$\frac{1}{p}\|(u_{n},v_{n})\|_{{\mathcal{W}}}^{p}-\frac{\gamma}{p_{s}^{*}}\int_{\Omega}{|u_{n}|^{\alpha}|v_{n}|^{\beta}}\,{\rm d}x\leq{\theta(\eta_{n})}+o_{n}(1).$
@@
$\left(\frac{1}{p}-\frac{1}{p_{s}^{*}}\right)\|(u_{n},v_{n})\|_{{\mathcal{W}}}^{p}\leq{\theta(\eta_{n})}+o_{n}(1)\leq\tilde{c}+o_{n}(1).$
@@
has a convergent subsequence, still denoted by $(\tilde{u}_{n},\tilde{v}_{n})$, such that $(\tilde{u}_{n},\tilde{v}_{n})\to(\tilde{u},\tilde{v})$ in ${\mathcal{W}}(\mathbb{R}^{d})$. Also, observe that
@@
Then we consider the following set
$\theta(\eta):=\inf_{{\mathcal{N}}_{\eta,\tau}^{\text{rad}}}I_{\eta,\tau}<\tilde{c}.$
Applying Lemma 5.2 and the Ekeland variational principle, there exists $(u_{\eta,\tau},v_{\eta,\tau})\in{\mathcal{N}}_{\eta,\tau}^{\text{rad}}$ such that $I_{\eta,\tau}(u_{\eta,\tau},v_{\eta,\tau})=\theta(\eta)$. Further, taking $\Omega=B_{\tau}$ in Lemma 4.8, we observe that
$\lim_{\eta\to 0}\theta(\eta)=\tilde{c}.$ (5.10)
@@
Now, we define the map $\zeta_{\tau}:\Omega_{\tau}^{-}\to I_{\eta}^{\theta(\eta)}$ by
@@
By either using spherical coordinates or the Divergence theorem, we can see that $\int_{\partial B_{\tau}}z_{i}\,{\rm d}S=0.$ Next, we define the map $H_{\eta}:[0,1]\times I_{\eta}^{\theta(\eta)}\to{\mathbb{R}}^{d}$ by
@@
There exists $\eta_{\ast\ast\ast}>0$ such that for $\eta\in(0,\eta_{\ast\ast\ast})$, $H_{\eta}\in{\mathcal{C}}([0,1]\times I^{\theta(\eta)},\Omega_{\tau}^{+}).$
@@
We first show that $H_{\eta}(t,(u,v))\in\Omega_{\tau}^{+}$ for $(u,v)\in I^{\theta(\eta)}$ and $t\in[0,1]$. On the contrary, suppose there exists $\eta_{n}\downarrow 0$, $t_{n}\to t_{0}\in[0,1]$ and $(u_{n},v_{n})\in I^{\theta(\eta_{n})}$ but $H_{\eta_{n}}(t_{n},(u_{n},v_{n}))\not\in\Omega_{\tau}^{+}$. By Lemma 5.5 and Lemma 5.4, we have
@@
$\text{cat}_{I_{\eta}^{\theta(\eta)}}(I_{\eta}^{\theta(\eta)})\geq\text{cat}_{\Omega}(\Omega).$
@@
Set $\eta_{\ast}:=\min\{\eta_{**},\eta_{***}\}$, where $\eta_{**}$ and $\eta_{***}$ are as in Lemma 5.4 and Lemma 5.6 respectively. Suppose $\text{cat}_{I_{\eta}^{\theta(\eta)}}(I_{\eta}^{\theta(\eta)})=N$. Then
$I_{\eta}^{\theta(\eta)}=F_{1}\cup\cdots\cup F_{N},$
where each $F_{i}$ is a closed and contractible set in $I_{\eta}^{\theta(\eta)}$. This means for each $1\leq i\leq N$, there exist $h_{i}\in{\mathcal{C}}([0,1]\times F_{i}:I_{\eta}^{\theta(\eta)})$ and tuples $(u_{i},v_{i})$ such that
@@
$\text{cat}_{I_{\eta}^{\theta(\eta)}}(I_{\eta}^{\theta(\eta)})\geq\text{cat}_{\Omega_{\tau}^{+}}(\Omega_{\tau}^{-})=\text{cat}_{\Omega}(\Omega),$

### v3

@@ 2. Existence and Non-existence of ground state solutions
Next, we prove the claims about $y_{n}$ and $r_{n}$. From (2.13) for large $n$,
$0<C_{3}=r_{n}^{d-sp}\fint_{B(y_{n},r_{n})}(|{u}_{n}|^{p}+|{v}_{n}|^{p})\,{\rm d}x\leq Cr_{n}^{-sp}\int_{\Omega}(|{u}_{n}|^{p}+|{v}_{n}|^{p})\,{\rm d}x.$
If $r_{n}\to\infty$, then using the fact that $\{u_{n}\},\{v_{n}\}$ are bounded in $L^{p}(\Omega)$, we get a contradiction. Thus, $\{r_{n}\}$ is a bounded sequence. Since $u_{n},v_{n}$ vanish in $\mathbb{R}^{d}\setminus\Omega$, using (2.13), we have $y_{n}\in\{x\in{\mathbb{R}}^{d}:d(x,\overline{\Omega})\leq r_{n}\}$. Therefore, $\{y_{n}\}$ is bounded in ${\mathbb{R}}^{d}$. Suppose, $r_{n}\to r_{0}>0$ and $y_{n}\to y\in{\mathbb{R}}^{d}$. Notice that $\tilde{u}_{n}$ and $\tilde{v}_{n}$ are supported on $\frac{\Omega-y_{n}}{r_{n}}$ and as $n\to\infty$,
@@
[…]
@@
[…]
@@ 3. Concentration compactness principle associated with the homogeneous system
Let $0\leq m\leq sp<d$ and $\alpha,\beta>1$ be such that $\alpha+\beta=p^{*}_{s}(m)$. Let $\{(u_{n},v_{n})\}$ weakly converges to (u,v){(u,v)} in $\mathcal{W}$. Then up to a subsequence, the following identities hold:
@@
Let $\Omega\subset{\mathbb{R}}^{d}$ be a bounded open set containing 0. Let $0\leq m\leq sp$ and $\alpha,\beta>1$ be such that $\alpha+\beta=p^{*}_{s}(m)$. Let $\{u_{n}\},\{v_{n}\}$ be sequences in $W_{0}^{s,p}(\Omega)$} such that $u_{n}\rightharpoonup u$ and $v_{n}\rightharpoonup v$ in $W_{0}^{s,p}(\Omega)$}. Then there exist measures $\sigma_{1},\sigma_{2},\nu_{1},\nu_{2},\nu$ and at most countable sets $\{x_{j}\}_{j\in\mathcal{J}_{1}}$, $\{x_{j}\}_{j\in\mathcal{J}_{2}}$, $\{x_{j}\}_{j\in\mathcal{J}_{3}}$, such that up to subsequences,
@@
which implies that $\{x_{j}\}_{j\in\mathcal{J}_{3}}=\{0\}$. Suppose there exists some $x_{j}\in\mathcal{J}_{3}$ with $x_{j}\neq 0$ and $x_{j}\in\text{supp}(\nu)$, then choosing $\phi\in{{\mathcal{C}}_{c}^{\infty}}(\Omega)$ and $\phi\equiv 1$ in a neighbourhood of $x_{j}$ with support does not contain 0, we get a contradiction repeating our previous calculation. Next, for $u\neq 0$ and $v\neq 0$, taking $\tilde{u}_{n}=u_{n}-u$ and $\tilde{v}_{n}=v_{n}-v$, we immediately get
@@ 4. Existence of a nontrivial solution with least energy
We say that $I$ satisfies PS condition at the level $c$ (in short, $\text{(PS)}_{c}$ condition), if $\{(u_{n},v_{n})\}$ is a PS sequence for $I$ at the level $c$ then $\{(u_{n},v_{n})\}$ has a convergent subsequence in $\mathcal{W}$}.
In the following lemma, depending on the values of $r$ and $m$, we see that the energy functional $I$ (defined in (1.4)) satisfies the $\text{(PS)}_{c}$ condition under a certain range of $c$.
Lemma 4.2 ($\text{(PS)}_{c}$ condition).
@@
    If $p_{s}^{*}(m)>p$ and $r\in\left(p,p_{s}^{*}\right)$, then for $\gamma,\eta>0$, $I$ satisfies $\text{(PS)}_{c}$ condition for every $c<\tilde{c}$.
@@
    If $p_{s}^{*}(m)>p$ and $r=p$ then for $\gamma>0$ and $\eta<\frac{S_{\alpha+\beta}}{c_{0}}$, $I$ satisfies $\text{(PS)}_{c}$ condition for every $c<\tilde{c}$.
@@
    If $p_{s}^{*}(m)=p$ and $r\in\left(p,p_{s}^{*}\right)$, then for $\gamma<S_{\alpha,\beta}$ and $\eta>0$, $I$ satisfies $\text{(PS)}_{c}$ condition for every $c\in\mathbb{R}$.
@@
Convergence of $\text{(PS)}_{c}$ sequence: Since $\mathcal{W}$ is reflexive, upto a subsequence, $\{(u_{n},v_{n})\}$ weakly converges to $(u,v)$ in $\mathcal{W}$. Hence $u_{n}\rightarrow u$ and $v_{n}\rightarrow v$ a.e. in $\Omega$. By [8, Lemma 2.5], we have the following convergence for every $(\phi,\psi)\in{\mathcal{W}}$:
@@
Assume $m<sp$, i.e. $p_{s}^{*}(m)>p$. By [27, Theorem 1.1], we choose a positive radially symmetric decreasing minimizer $U_{m}=U_{m}(r)$ for $S_{\alpha+\beta}$ which weakly solves
@@
For $\varepsilon>0$, the function $U_{m,\varepsilon}(x)=\varepsilon^{-\frac{N-ps}{p}}U_{m}\left(\frac{x}{\varepsilon}\right)$ is also a minimizer of $S_{\alpha+\beta}$ satisfying (4.11). Now, for $\delta>0$, we define the radially symmetric non-negative decreasing function
@@
[…]
@@
$I\left(tu_{0},tv_{0}\right)=g_{\varepsilon}(t)$ $=\frac{1}{p}\left(\left[tu_{0}\right]_{s,p}^{p}+\left[tv_{0}\right]_{s,p}^{p}\right)-\frac{\gamma}{p_{s}^{*}(m)}\int_{\Omega}\frac{\left(tu_{0}\right)^{\alpha}\left(tv_{0}\right)^{\beta}}{|x|^{m}}\,{\rm d}x$
@@
$\quad\quad-\frac{\eta t^{r}}{r}\left(\alpha^{\frac{r}{p}}+\beta^{\frac{r}{p}}\right)\int_{\Omega}u_{m,\varepsilon,1}^{r}\,{\rm d}x$
@@
$I(u~,v~),l_{S}:=\displaystyle\inf_{(u,v)\in S}$$I(\tilde{u},\tilde{v})$,
@@
[…]
@@
[…]
@@
$u]_{s,p}^{p}+[v]_{s,p}^{p}=t^{p_{s}^{\ast}(m)-p}\gamma\int_{\Omega}\frac{\left|u\right|^{\alpha}\left|v\right|^{\beta}}{|x|^{m}}\,{\rm d}x+t^{r-p}\eta\int_{\Omega}\left(\left|u\right|^{r}+\left|v\right|^{r}\right)\,{\rm d}x.$
@@
[…]
$\theta(\tilde{\eta}):=\inf_{{\mathcal{N}}_{\tilde{\eta}}}I_{\tilde{\eta}}$.
@@
[…]
@@ 5. Existence of category many nontrivial solutions
[…]
where $c_{0}$ is given in (1.5). We define $\psi_{\eta}:{\mathcal{W}}\to{\mathbb{R}}$ by
@@
We say that $I_{\eta}$ restricted on ${\mathcal{N}}_{\eta}$ satisfies PS condition at level $c$, $\text{(PS)}_{c}^{{\mathcal{N}}_{\eta}}$ in short, if for $\{(u_{n},v_{n})\}\subset{\mathcal{N}}_{\eta}$ such that
@@
Then we consider the following set
$\theta(\eta)\leq\theta_{1}(\eta):=\inf_{{\mathcal{N}}_{\eta,\tau}^{\text{rad}}}I_{\eta,\tau}<\tilde{c}$.
Applying Lemma 5.2 and the Ekeland variational principle, there exists $(u_{\eta,\tau},v_{\eta,\tau})\in{\mathcal{N}}_{\eta,\tau}^{\text{rad}}$ such that $I_{\eta,\tau}(u_{\eta,\tau},v_{\eta,\tau})=\theta_{1}(\eta)$. Now our aim is to show that for $\eta$ small enough, $\text{cat}_{I_{\eta}^{\theta_{1}(\eta)}}(I_{\eta}^{\theta_{1}(\eta)})$ is bigger than $\text{cat}_{\Omega}(\Omega).$ For $\tau>0$, we define
@@
There exists $\eta_{\ast\ast}>0$ such that for $\eta\in(0,\eta_{\ast\ast})$, $\Phi_{\eta}(u,v)\in\Omega_{\tau}^{+}$ for every $(u,v)\in I_{\eta}^{\theta_{1}(\eta)}$.
@@
Arguing by contradiction, suppose there exists $\eta_{n}>0$, $\eta_{n}\to 0$ and $(u_{n},v_{n})\in I_{\eta_{n}}^{\theta_{1}(\eta_{n})}$ such that $\Phi_{\eta_{n}}(u_{n},v_{n})\not\in\Omega_{\tau}^{+}$. Then
@@
$\frac{1}{p}\|(u_{n},v_{n})\|_{{\mathcal{W}}}^{p}-\frac{\gamma}{p_{s}^{*}}\int_{\Omega}{|u_{n}|^{\alpha}|v_{n}|^{\beta}}\,{\rm d}x-\frac{\eta_{n}}{r}\int_{\Omega}(|u_{n}|^{r}+|v_{n}|^{r})\,{\rm d}x\leq{\theta_{1}(\eta_{n})}<\tilde{c}$. (5.4)
@@
$\frac{1}{p}\|(u_{n},v_{n})\|_{{\mathcal{W}}}^{p}-\frac{\gamma}{p_{s}^{*}}\int_{\Omega}{|u_{n}|^{\alpha}|v_{n}|^{\beta}}\,{\rm d}x\leq{\theta_{1}(\eta_{n})}+o_{n}(1).$
@@
$\left(\frac{1}{p}-\frac{1}{p_{s}^{*}}\right)\|(u_{n},v_{n})\|_{{\mathcal{W}}}^{p}\leq{\theta_{1}(\eta_{n})}+o_{n}(1)\leq\tilde{c}+o_{n}(1).$
@@
[…]
@@
Taking $\Omega=B_{\tau}$ in Lemma 4.8, we observe that
$\lim_{\eta\to 0}\theta_{1}(\eta)=\tilde{c}$. (5.10)
@@
Now, we define the map $\zeta_{\tau}:\Omega_{\tau}^{-}\to I_{\eta}^{\theta_{1}(\eta)}$ by
@@
By either using spherical coordinates or the Divergence theorem, we can see that $\int_{\partial B_{\tau}}z_{i}\,{\rm d}S=0.$ Next, we define the map $H_{\eta}:[0,1]\times I_{\eta}^{\theta_{1}(\eta)}\to{\mathbb{R}}^{d}$ by
@@
There exists $\eta_{\ast\ast\ast}>0$ such that for $\eta\in(0,\eta_{\ast\ast\ast})$, $H_{\eta}\in{\mathcal{C}}([0,1]\times I^{\theta_{1}(\eta)},\Omega_{\tau}^{+}).$
@@
We first show that $H_{\eta}(t,(u,v))\in\Omega_{\tau}^{+}$ for $(u,v)\in I^{\theta_{1}(\eta)}$ and $t\in[0,1]$. On the contrary, suppose there exists $\eta_{n}\downarrow 0$, $t_{n}\to t_{0}\in[0,1]$ and $(u_{n},v_{n})\in I^{\theta_{1}(\eta_{n})}$ but $H_{\eta_{n}}(t_{n},(u_{n},v_{n}))\not\in\Omega_{\tau}^{+}$. By Lemma 5.5 and Lemma 5.4, we have
@@
$\text{cat}_{I_{\eta}^{\theta_{1}(\eta)}}(I_{\eta}^{\theta_{1}(\eta)})\geq\text{cat}_{\Omega}(\Omega).$
@@
Set $\eta_{\ast}:=\min\{\eta_{**},\eta_{***}\}$, where $\eta_{**}$ and $\eta_{***}$ are as in Lemma 5.4 and Lemma 5.6 respectively. Suppose $\text{cat}_{I_{\eta}^{\theta_{1}(\eta)}}(I_{\eta}^{\theta_{1}(\eta)})=N$. Then
$I_{\eta}^{\theta_{1}(\eta)}=F_{1}\cup\cdots\cup F_{N},$
where each $F_{i}$ is a closed and contractible set in $I_{\eta}^{\theta_{1}(\eta)}$. This means for each $1\leq i\leq N$, there exist $h_{i}\in{\mathcal{C}}([0,1]\times F_{i}:I_{\eta}^{\theta_{1}(\eta)})$ and tuples $(u_{i},v_{i})$ such that
@@
$\text{cat}_{I_{\eta}^{\theta_{1}(\eta)}}(I_{\eta}^{\theta_{1}(\eta)})\geq\text{cat}_{\Omega_{\tau}^{+}}(\Omega_{\tau}^{-})=\text{cat}_{\Omega}(\Omega),$